# A restless Time-Fractional Multiclass Queue

### v1

@@ 1. Introduction and preliminaries
The theory of queues gives foundational mathematical tools for rigorous modelling of congestion across different disciplines including computing, communications, logistics, and healthcare. Classical models such as M/M/1 and their GI/GI/1 generalisations are analytically tractable and offer insight, as well as genuine modelling methodologies for practitioners, through a vast body of literature [12, 5, 29]. In the studies of such models, the common assumptions are the memoryless Markovian properties or the light-tailness of distributions of the service times or arrival times. The performance and stability of such queues are essentially governed by the moments of these times. Of a particular importance is the notion of ‘traffic intensity’, through which the stability properties of the queue such as recurrence, transience, laws of large numbers or fluctuations of the queue length are completely characterised.
@@
This creates a need to develop a theory for queues in which the distributions of services or arrivals (or both) create long-range effects. One way to construct such queues is to enforce the inter-arrival or service times to be heavy-tailed, in the sense of having infinite mean. This behaviour has already been observed in examples [28, 30].
alter the behaviour of all quantities of interest in the queue, such as system delays, tail asymptotics for waiting times, order of queue lengths, and busy periods as the process evolves in time [20, 26, 16]. In such queueing examples the well-trodden traffic intensity approach loses its meaning and one must characterise regimes via scaling limits and reflected-input structures rather than expectations of arrival and service times [1, 10].
@@
  1. Model 1:
    At the event times of a standard Poisson process, the queue length (i) increases by $1$ with probability $1/2$, or (ii) with probability $1/2$, either decreases by $1$ if at least one customer is present, or stays 0 if there are no customers.
  2. Model 2:
    Consider two independent sequences of independent Exp(1/2) times $\{A_{i}\}_{i\geq 1}$ and $\{B_{i}\}_{i\geq 1}$. Consider the sequence of independent times $T_{i}=A_{i}\wedge B_{i}\sim{\rm Exp}(1)$ to form a standard Poisson process with the inter-event times $T_{i}$. The queue length changes at these times so that a customer arrives (+1) at the $i$-th event if $A_{i}<B_{i}$, otherwise a customer leaves (1-1) or the queue stays empty, as above.
  3. Model 3:
    Use two independent Poisson(1/2) processes $N_{a}$ and $N_{d}$. Whenever $N_{a}$ has an event, then a customer arrives. Whenever $N_{d}$ has an event, a customer leaves (1-1) or the queue stays empty.
  4. Model 4:
    Use one Poisson(1/2) process $N_{a}$ for arrivals. The first customer in the queue exits after an independent $\rm Exp(1)$ service time passes. If the queue is empty, no service times run.
@@ 4. Model 4:
In the current article we focus on a multi-class queueing model where both arrivals and service completions are events times of two independent fractional Poisson processes, which makes this the multiclass generalisation of fractional Model 3 above.
@@
Acknowledgments: NG and VV acknowledge the support of the Dr Perry James (Jim) Browne Research Center at the Department of Mathematics, University of Sussex. ES was partially supported by the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 published on 2.2.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU– Project Title “Non–Markovian Dynamics and Non-local Equations” – 202277N5H9 - CUP: D53D23005670006 - Grant Assignment Decree No. 973 adopted on June 30, 2023, by the Italian Ministry of University and Research (MUR) and by the project “Approximation of stochastic processes by means of sums of random telegraph processes” funded by Sapienza University of Rome. NG would also like to thank Giacomo Ascione for discussions on fractional queueing systems and for early access to the preprint [1].
@@ 1.3. Time change
The random variable $Y_{\theta}(1)$ has density when $\theta\in(0,1)$, which we denote by $h_{\theta}(x)$; by the self-similarity, $Y_{\theta}(t)$ has densitiy $t^{-\theta}h_{\theta}(xt^{-\theta})$ when $t>0$. For $\theta=1$, we have $Y_{1}(t)=t$ for $t\geq 0$.
@@ 2.1. The thinned FPP
(2.4) ${\rm Cov}(N^{(i)}(t),N^{(j)}(t))=p_{i}p_{j}\lambda^{2\alpha}t^{2\alpha}\left(\frac{1}{\Gamma(1+2\alpha)}-\frac{1}{(\Gamma(1+\alpha))^{2}}\right)=p_{i}p_{j}\lambda^{2\alpha}{\rm Var}(Y_{\alpha}(t)).$
@@
${\rm Cov}(N^{(i)}(t),N^{(j)}(t))|N^{\alpha,\lambda}(t))$ $=\mathbb{E}(N^{(i)}(t)N^{(j)}(t)|N^{\alpha,\lambda}(t))-\mathbb{E}(N^{(i)}(t)|N^{\alpha,\lambda}(t))\mathbb{E}(N^{(j)}(t)|N^{\alpha,\lambda}(t))$
@@ 2.2. The discrete-class multiclass queue: Scaling limits and recurrence
Assume, as in Theorem 2.7, that the class arrival process is given by $(N_{1}(p_{1}Y_{\alpha}(t)),\ldots,N_{K}(p_{K}Y_{\alpha}(t)))$, and the service departure process is $\widetilde{N}(\widetilde{Y}_{\beta}(t))$, where $N_{1},\ldots,N_{K}$ are Poisson processes of rate $\lambda^{\alpha}$ and $\widetilde{N}$ is a Poisson process of rate $\mu^{\beta}$, $Y_{\alpha}$ and $Y_{\beta}$ are inverse standard stable subordinators of respective indices $\alpha$ and $\beta$, and all the $K+3$ processes are independent. For fixed $i,\lambda,\mu,p_{1},\ldots,p_{i}$, which will not appear in the shorthanded notation below, define
@@
We believe that $H_{\alpha,\beta}(u\cdot)/u^{\gamma/2}$ has a non-trivial limit, and as such we should be able to obtain scaling limit for $\widetilde{Q}^{\rm cent}_{\leq i}$ This study will be left as part of a future work.
@@
When $\alpha<\beta$, the (total) arrival process is slower than the departure process, so the queue will empty infinitely often with probability 1. This is Theorem 3.4 in [10]. The same theorem gives that if $\alpha>\beta$, then $\limsup$ of the queue length is $\infty$ with probability 1. Moreover, by Theorem 3.3 in [10], the scaling limit of the queue length would be distributionally equal to the inverse $\alpha$-stable subordinator. Here, both theorems can be applied to the queue length of Class 1 since the arrival process for it is FPP by Theorem 2.1, to show corresponding results.
@@ 3. Proofs for the arrival process
(3.4) $\big({\bf S}(N(y_{1})),\ldots,{\bf S}(N(y_{n}))\big)\stackrel{{\scriptstyle d}}{{=}}\Big(\big(N_{1}(p_{1}y_{1}),\ldots,N_{K}(p_{K}y_{1})\big),\ldots,\big(N_{1}(p_{1}y_{n}),\ldots,N_{K}(p_{K}y_{n})\big)\Big).$
This is equivalent [^1] Equalities (3.4) all together mean that $\{{\bf S}(N(y)\}_{y\geq 0}\stackrel{{\scriptstyle d}}{{=}}\big\{\big(N_{1}(p_{1}y),\ldots,N_{K}(p_{K}y)\big)\big\}_{y\geq 0},$ which describes the classical thinning of the Poisson processes $N$ using $K$ marks with probabilities $p_{i}$. Indeed, the l.h.s. is simply the multivariate process of class counts in our queue if we assume that the process of arrivals is $N$. The marginals of ${\bf S}(N(y))$ are mutually independent Poisson processes of rate $p_{i}\lambda^{\alpha}$, according to the thinning. However, we prefer to give a full proof of  (3.4). to establishing the following joint distributional identity for the increments:
@@
$\stackrel{{\scriptstyle d}}{{=}}\Big(\big(N_{1}(p_{1}y_{1}),\ldots,N_{K}(p_{K}y_{1})\big),\ldots,\big(N_{1}(p_{1}y_{n})-N_{1}(p_{1}y_{n-1}),\ldots,N_{K}(p_{K}y_{n})-N_{K}(p_{K}y_{n-1})\big)\Big).$
@@
${\rm Cov}(N^{(i)}(t),N^{(j)}(t))$ $=\mathbb{E}({\rm Cov}(N^{(i)}(t),N^{(j)}(t))|N^{\alpha,\lambda}(t))+{\rm Cov}(E(N^{(i)}(t)|N^{\alpha,\lambda}(t)),E(N^{(j)}(t)|N^{\alpha,\lambda}(t)))$
@@
$=p_{i}p_{j}\big[\mathbb{E}({\rm Var}(N^{\alpha,\lambda}(t)|Y_{\alpha}))+{\rm Var}(E(N^{\alpha,\lambda}(t)|Y_{\alpha}(t)))-\mathbb{E}(E(N^{\alpha,\lambda}(t)|Y_{\alpha}(t)))\big]$
@@
$=\mathbb{P}\Big(N^{1,\lambda^{\alpha}P_{\ell}}(y_{1})=k_{1},\ldots,N^{1,\lambda^{\alpha}P_{\ell}}(y_{n})=k_{n}\Big)$
@@
$\sup_{0\leq t\leq T}\Big|\tilde{\bf N}_{u}(t)$ $-\lambda^{\alpha}u^{-\alpha}{\bf p}Y_{\alpha}(ut)\Big|\leq|\tilde{\bf N}_{u}(1/\sqrt{u})|+\lambda^{\alpha}u^{-\alpha}Y_{\alpha}(\sqrt{u})|{\bf p}|+\sup_{1/\sqrt{u}\leq t\leq T}\Big|\tilde{\bf N}_{u}(t)-\lambda^{\alpha}u^{-\alpha}{\bf p}Y_{\alpha}(ut)\Big|$
(3.7) $\leq u^{-\alpha}Y_{\alpha}(\sqrt{u})|L(\sqrt{u})|+\lambda^{\alpha}u^{-\alpha}Y_{\alpha}(\sqrt{u})|{\bf p}|+u^{-\alpha}Y_{\alpha}(uT)\sup_{1/\sqrt{u}\leq t\leq T}\big|L(ut)-\lambda^{\alpha}{\bf p}\big|.$
By the strong law of large numbers applied to the Poisson processes $N_{1},\ldots,N_{K}$ and the fact that $Y_{\alpha}(t)\to\infty$ a.s. as $t\to\infty$, we have $L(t)\to\lambda^{\alpha}{\bf p}$ a.s. as $t\to\infty$; to see this, divide and multiply by $p_{i}$ each coordinate $i$ of $L(t)$. Therefore, the last supremum in (3) vanishes a.s. as $u\to\infty$. Combined with the distributional equalities $u^{-\alpha}Y_{\alpha}(\sqrt{u})\stackrel{{\scriptstyle d}}{{=}}u^{-\alpha/2}Y_{\alpha}(1)$ and $u^{-\alpha}Y_{\alpha}(uT)\stackrel{{\scriptstyle d}}{{=}}Y_{\alpha}(T)$, this implies that the r.h.s. of (3) converges to 0 in probability, establishing (3.6).
@@ 4. Proofs for the queueing system
In this section we prove the results for the restless multiclass queue. In Subsection 4.1 we deal with the first-order scaling limits in the spirit of Theorem 3.3 in [10] and an FCLT as corollary of a compensated queue length. In Subsection 4.2, we prove that the restless (multiclass) queue empties infinitely often when the indices of the arrival and service FPPs are equal.
@@ 4.1. Scaling limits for queue lengths
After that the first claim in the theorem follows, since the infimum process only decreases by 1 at the times when a service was completed but the queue was empty, by (1.15) and Lemma 6.2. in [10].
@@
The sequences $\{W_{n}\}_{n}$ and $\{\widetilde{W}_{n}\}_{n}$ are i.i.d. respectively as increments of Lévy processes over fixed size intervals, and jointly independent. This makes $S_{n}$ a random walk whose i.i.d. increments $J_{n}$ are strictly stable and their common distribution puts mass on both half-axes $(-\infty,0)$ and $(0,\infty)$, that is $\mathbb{P}(J_{n}>0)>0$ and $\mathbb{P}(J_{n}<0)>0$.

### v2

@@ 1. Introduction and preliminaries
The theory of queues gives foundational mathematical tools for rigorous modelling of congestion across different disciplines including computing, communications, logistics, and healthcare. Classical models such as M/M/1 and their GI/GI/1 generalisations are analytically tractable and offer insight, as well as genuine modelling methodologies for practitioners, through a vast body of literature [12, 5, 29]. In the studies of such models, the common assumptions are the memoryless Markovian properties or the light-tailness of distributions of the service times or arrival times. The performance and stability of such queues are essentially governed by the moments of these times. Of particular importance is the notion of ‘traffic intensity’, through which the stability properties of the queue such as recurrence, transience, laws of large numbers or fluctuations of the queue length are completely characterised.
@@
This creates a need to develop a theory for queues in which the distributions of services or arrivals (or both) create long-range effects. One way to construct such queues is to enforce the inter-arrival or service times to be heavy-tailed, in the sense of having infinite mean. This behaviour has already been observed in examples [28, 30]. Such changes alter the behaviour of all quantities of interest in the queue, such as system delays, tail asymptotics for waiting times, order of queue lengths, and busy periods as the process evolves in time [20, 26, 16]. In such queueing examples the well-trodden traffic intensity approach loses its meaning and one must characterise regimes via scaling limits and reflected-input structures rather than expectations of arrival and service times [1, 10].
@@
Model 1: At the event times of a standard Poisson process, the queue length (i) increases by $1$ with probability $1/2$, or (ii) with probability $1/2$, either decreases by $1$ if at least one customer is present, or stays 0 if there are no customers.
Model 2: Consider two independent sequences of independent Exp(1/2) times $\{A_{i}\}_{i\geq 1}$ and $\{B_{i}\}_{i\geq 1}$. Consider the sequence of independent times $T_{i}=A_{i}\wedge B_{i}\sim{\rm Exp}(1)$ to form a standard Poisson process with the inter-event times $T_{i}$. The queue length changes at these times so that a customer arrives (+1) at the $i$-th event if $A_{i}<B_{i}$, otherwise a customer leaves (1-1) or the queue stays empty, as above.
Model 3: Use two independent Poisson(1/2) processes $N_{a}$ and $N_{d}$. Whenever $N_{a}$ has an event, then a customer arrives. Whenever $N_{d}$ has an event, a customer leaves (1-1) or the queue stays empty.
Model 4: Use one Poisson(1/2) process $N_{a}$ for arrivals. The first customer in the queue exits after an independent $\rm Exp(1)$ service time passes. If the queue is empty, no service times run.
@@
In the current article we focus on a multi-class queueing model where both arrivals and service completions are event times of two independent fractional Poisson processes, which makes this the multiclass generalisation of fractional Model 3 above.
@@
Acknowledgments: NG and VV acknowledge the support of the Dr Perry James (Jim) Browne Research Center at the Department of Mathematics, University of Sussex. NG acknowledges financial support provided by Sapienza University of Rome through the program Professori Visitatori 2025.
ES was partially supported by the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 published on 2.2.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU– Project Title “Non–Markovian Dynamics and Non-local Equations” – 202277N5H9 - CUP: D53D23005670006 - Grant Assignment Decree No. 973 adopted on June 30, 2023, by the Italian Ministry of University and Research (MUR). ES also acknowledges financial support provided by Sapienza University of Rome 000317 24 RICERCA UNIV 2023 PROG MEDI SCALAS - RICERCA ATENEO 2023 - SCALAS PROGETTI MEDI. The title of the project is “Approximation of stochastic processes by means of sums of random telegraph processes”.
The authors would like to thank the anonymous referees for their comments and suggestions. NG would also like to thank Giacomo Ascione for discussions on fractional queueing systems and for early access to the preprint [1].
@@ 1.3. Time change
The random variable $Y_{\theta}(1)$ has density when $\theta\in(0,1)$, which we denote by $h_{\theta}(x)$; by the self-similarity, $Y_{\theta}(t)$ has density $t^{-\theta}h_{\theta}(xt^{-\theta})$ when $t>0$. For $\theta=1$, we have $Y_{1}(t)=t$ for $t\geq 0$.
@@ 2.1. The thinned FPP
(2.4) ${\rm Cov}(N^{(i)}(t),N^{(j)}(t))$ $=p_{i}p_{j}\lambda^{2\alpha}t^{2\alpha}\left(\frac{1}{\Gamma(1+2\alpha)}-\frac{1}{(\Gamma(1+\alpha))^{2}}\right)$
(2.5) $=p_{i}p_{j}\lambda^{2\alpha}{\rm Var}(Y_{\alpha}(t)).$
@@
${\rm Cov}(N^{(i)}(t),N^{(j)}(t))|N^{\alpha,\lambda}(t))$ $=\mathbb{E}(N^{(i)}(t)N^{(j)}(t)|N^{\alpha,\lambda}(t))$
$\phantom{xxxxxxxxxxxx}-\mathbb{E}(N^{(i)}(t)|N^{\alpha,\lambda}(t))\mathbb{E}(N^{(j)}(t)|N^{\alpha,\lambda}(t))$
@@ 2.2. The discrete-class multiclass queue: Scaling limits and recurrence
Assume, as in Theorem 2.7, that the class arrival process is given by $(N_{1}(p_{1}Y_{\alpha}(t)),\ldots,N_{K}(p_{K}Y_{\alpha}(t)))$, and the service departure process is $\widetilde{N}(\widetilde{Y}_{\beta}(t))$, where $N_{1},\ldots,N_{K}$ are Poisson processes of rate $\lambda^{\alpha}$ and $\widetilde{N}$ is a Poisson process of rate $\mu^{\beta}$, $Y_{\alpha}$ and $\widetilde{Y}_{\beta}$ are inverse standard stable subordinators of respective indices $\alpha$ and $\beta$, and all the $K+3$ processes are independent. For fixed $i,\lambda,\mu,p_{1},\ldots,p_{i}$, which will not appear in the shorthanded notation below, define
@@
We believe that $H_{\alpha,\beta}(u\cdot)/u^{\gamma/2}$ has a non-trivial limit, and as such we should be able to obtain scaling limit for $\widetilde{Q}^{\rm cent}_{\leq i}$. This study will be left as part of a future work.
@@
When $\alpha<\beta$, the (total) arrival process is slower than the departure process, so the queue will empty infinitely often with probability 1. This is Theorem 3.4 in [10] (see also Theorem A.2 in Appendix A). The same theorem gives that if $\alpha>\beta$, then $\limsup$ of the queue length is $\infty$ with probability 1. Moreover, by Theorem 3.3 in [10] (Theorem A.1 in Appendix A), the scaling limit of the queue length would be distributionally equal to the inverse $\alpha$-stable subordinator. Here, both theorems can be applied to the queue length of Class 1 since the arrival process for it is FPP by Theorem 2.1, to show corresponding results.
For completion we record the statements, but it is a direct application of Theorem 3.4. in [10] (first part of Theorem A.2 in Appendix A so the proof will be omitted.
Theorem 2.14.
The following statements hold true, irrespective of the values of $\lambda,\mu,p_{i}$. For any fixed $i\leq K$,
  1. (1)
    If $\alpha<\beta$, the queue empties infinitely often with probability $1$.
  2. (2)
[…]
@@ 3. Proofs for the arrival process
$\big({\bf S}(N(y_{1})),\ldots,$ ${\bf S}(N(y_{n}))\big)$
(3.4) $\stackrel{{\scriptstyle d}}{{=}}\Big(\big(N_{1}(p_{1}y_{1}),\ldots,N_{K}(p_{K}y_{1})\big),\ldots,\big(N_{1}(p_{1}y_{n}),\ldots,N_{K}(p_{K}y_{n})\big)\Big).$
This is equivalent  [^1] Equalities (3) all together mean that $\{{\bf S}(N(y)\}_{y\geq 0}\stackrel{{\scriptstyle d}}{{=}}\big\{\big(N_{1}(p_{1}y),\ldots,N_{K}(p_{K}y)\big)\big\}_{y\geq 0},$ which describes the classical thinning of the Poisson processes $N$ using $K$ marks with probabilities $p_{i}$. Indeed, the l.h.s. is simply the multivariate process of class counts in our queue if we assume that the process of arrivals is $N$. The marginals of ${\bf S}(N(y))$ are mutually independent Poisson processes of rate $p_{i}\lambda^{\alpha}$, according to the thinning. However, we prefer to give a full proof of  (3). to establishing the following joint distributional identity for the increments:
@@
$\stackrel{{\scriptstyle d}}{{=}}\Big(\big(N_{1}(p_{1}y_{1}),\ldots,N_{K}(p_{K}y_{1})\big),\ldots,$
$\phantom{xxxxxxxx}\big(N_{1}(p_{1}y_{n})-N_{1}(p_{1}y_{n-1}),\ldots,N_{K}(p_{K}y_{n})-N_{K}(p_{K}y_{n-1})\big)\Big).$
@@
${\rm Cov}(N^{(i)}(t),N^{(j)}(t))$ $=\mathbb{E}({\rm Cov}(N^{(i)}(t),N^{(j)}(t))|N^{\alpha,\lambda}(t))$
$\phantom{xxxxxxx}+{\rm Cov}(E(N^{(i)}(t)|N^{\alpha,\lambda}(t)),E(N^{(j)}(t)|N^{\alpha,\lambda}(t)))$
@@
$=p_{i}p_{j}\big[\mathbb{E}({\rm Var}(N^{\alpha,\lambda}(t)|Y_{\alpha}))$
$\phantom{xxxxxxx}+{\rm Var}(E(N^{\alpha,\lambda}(t)|Y_{\alpha}(t)))-\mathbb{E}(E(N^{\alpha,\lambda}(t)|Y_{\alpha}(t)))\big]$
@@
$=\mathbb{P}\Big(N^{1,\lambda^{\alpha}P_{\ell}}(y_{1})=k_{1},\ldots,N^{1,\lambda^{\alpha}P_{\ell}}(y_{n})=k_{n}\Big).$
@@
$\sup_{0\leq t\leq T}\Big|\tilde{\bf N}_{u}(t)$ $-\lambda^{\alpha}u^{-\alpha}{\bf p}Y_{\alpha}(ut)\Big|$
$\leq|\tilde{\bf N}_{u}(1/\sqrt{u})|+\lambda^{\alpha}u^{-\alpha}Y_{\alpha}(\sqrt{u})|{\bf p}|+\sup_{1/\sqrt{u}\leq t\leq T}\Big|\tilde{\bf N}_{u}(t)-\lambda^{\alpha}u^{-\alpha}{\bf p}Y_{\alpha}(ut)\Big|$
$\leq u^{-\alpha}Y_{\alpha}(\sqrt{u})|L(\sqrt{u})|+\lambda^{\alpha}u^{-\alpha}Y_{\alpha}(\sqrt{u})|{\bf p}|$
(3.7) $\phantom{xxxxxxxxxxxxxxxxxx}\!+u^{-\alpha}Y_{\alpha}(uT)\sup_{1/\sqrt{u}\leq t\leq T}\big|L(ut)-\lambda^{\alpha}{\bf p}\big|.$
By the strong law of large numbers applied to the Poisson processes $N_{1},\ldots,N_{K}$ and the fact that $Y_{\alpha}(t)\to\infty$ a.s. as $t\to\infty$, we have $L(t)\to\lambda^{\alpha}{\bf p}$ a.s. as $t\to\infty$; to see this, divide and multiply by $p_{i}$ each coordinate $i$ of $L(t)$. Therefore, the last supremum in (3.7) vanishes a.s. as $u\to\infty$. Combined with the distributional equalities $u^{-\alpha}Y_{\alpha}(\sqrt{u})\stackrel{{\scriptstyle d}}{{=}}u^{-\alpha/2}Y_{\alpha}(1)$ and $u^{-\alpha}Y_{\alpha}(uT)\stackrel{{\scriptstyle d}}{{=}}Y_{\alpha}(T)$, this implies that the r.h.s. of (3.7) converges to 0 in probability, establishing (3.6).
@@ 4. Proofs for the queueing system
In this section we prove the results for the restless multiclass queue. In Subsection 4.1 we deal with the first-order scaling limits in the spirit of Theorem 3.3 in [10] (see for convenience Appendix A, theorem A.1) and an FCLT as corollary of a compensated queue length. In Subsection 4.2, we prove that the restless (multiclass) queue empties infinitely often when the indices of the arrival and service FPPs are equal.
@@ 4.1. Scaling limits for queue lengths
After that the first claim in the theorem follows, since the infimum process only decreases by 1 at the times when a service was completed but the queue was already empty, by (1.15) and Lemma 6.2. in [10] (see Theorem A.4 in Appendix A).
@@
The sequences $\{W_{n}\}_{n}$ and $\{\widetilde{W}_{n}\}_{n}$ are i.i.d. respectively as increments of Lévy processes over fixed size intervals, and jointly independent. These random variables are strictly $\alpha$-stable, and therefore for any $a,b>0$,
$aJ_{1}+bJ_{2}=(aW_{1}+bW_{2})+(a\widetilde{W}_{1}+b\widetilde{W}_{2})\stackrel{{\scriptstyle d}}{{=}}(a^{\alpha}+b^{\alpha})^{1/\alpha}W_{1}+(a^{\alpha}+b^{\alpha})^{1/\alpha}\widetilde{W}_{1}=(a^{\alpha}+b^{\alpha})^{1/\alpha}J_{1}.$
This makes $S_{n}$ a random walk whose i.i.d. increments $J_{n}$ are strictly stable and their common distribution puts mass on both half-axes $(-\infty,0)$ and $(0,\infty)$, that is $\mathbb{P}(J_{n}>0)>0$ and $\mathbb{P}(J_{n}<0)>0$.